# Observational predictions of some inflationary models

### v3

@@ Abstract
This paper presents the CMB angular power spectrum obtained using the CAMB code for three different models of inflation: the Starobinsky inflationary model, the generalized Starobinsky inflationary model, and the chaotic inflationary model with a step. The results are compared with the most recent data reported for the Planck mission. An analysis of the large ($l\lesssim 90$), intermediate ($90\lesssim l\lesssim 900$), and small ($l\gtrsim 900$) angular scales is performed. We report the position of the peaks in the intermediate region so as the cosmological parameters obtained in each of the models: age of the universe, $\Omega_{m}$, $\Omega_{b}$, $\Omega_{\Lambda}$, $\Omega_{K}$ and $n_{\mathrm{S}}$.
@@
CMB angular power spectrum , CAMB code , Starobinsky inflationary model , generalized Starobinsky inflationary model , chaotic inflationary model with a step.
@@ 1 Introduction
The Big Bang theory describes the development of the Universe from a much hotter and denser point to the present. It is supported by solid observational evidence: the expansion of the Universe, the prediction of light element abundances, and the existence of the Cosmic Microwave Background (CMB) radiation. The CMB is a picture of the recombination epoch ($370,000$ years after the big bang). The first prediction of the CMB was made by Alpher and Herman in $1948$ [1], and was discovered in $1965$ by Penzias and Wilson [2]. It became one of the most important observational probes of the Big–bang theory. The CMB is an almost uniform and isotropic radiation field and shows a perfect black body spectrum at a temperature of $2.72$K [3]. The CMB has fluctuations in a part of $10^{5}$ that give rise to the structure of the present Universe [4].
Although the big bang theory stands as a robust and widely accepted explanation for the beginning of our Universe, this theory has three problems related to its initial conditions, the flatness problem, the horizon problem, and the monopole problem. To address these concerns, Alan Guth introduced the inflationary theory in $1981$ [5]. This theory proposes an epoch of accelerated expansion preceding the hot stage mentioned in the traditional Big Bang model. The simplest possibility to produce such epoch is via the potential energy of a scalar field, and the field $\phi$ is called the inflaton. There are several inflationary models with their own scalar potentials, and through the years studies on which of them are favored by the satellites data have been made [6, 7, 8, 9]. One notable outcome of inflation is that provides a mechanism to generate the primordial scalar perturbations and produce an almost scale invariant scalar power spectrum.
@@
Using the CAMB code we study for each model the three regions of the CMB angular power spectrum: a) the Sachs–Wolfe plateau region ($\ell<90$) which correspond to large angular scales ($\theta>2^{\circ}$), b) the acoustic peak region ($90090<\ell<$ 900), and c) the silk damping region ($\ell>900$). It is important to note that Macorra et al. have used the CAMB code with full COSMO--MC (MCMC) calculation tzo give a solution to one of the most recent conflicts in Modern Cosmology: the Hubble tension [23]. Also, Purkayastha et al. have estimate CMB E mode signal over large angular scales [24].
To study CMB, we utilize the radiation angular power spectrum, which is both measurable and theoretically predictable. The radiation angular power spectrum $C_{\ell}$ is defined by:
@@
$C_{\ell}$ depends only of $\ell$ since the statistical properties are required to be independent of the choice of the origin of $\theta-\phi$, that is rotational invariance, or from a physical point of view, isotropy [25]. The larger is $\ell$, the smaller is the angular scale at which the spherical harmonics have variation [26].
Also (and it is the case for this article), the $C_{\ell}$ spectrum is conventionally plotted as $\ell(\ell+1)C_{\ell}\overline{T}^{2}/{2\pi}$, measuring the power per logarithmic interval in $\ell$. In such plot, a scale–invariant spectrum looks horizontal [25].
@@ 1.1 The Sachs-–Wolfe plateau region ($\ell$ $<$ 90)
Assuming a nearly scale–invariant spectrum of density perturbations in early times ($n_{S}\approx 1$), meaning gravitational potential fluctuations that are independent of $k$, then $\ell(\ell+1)C_{\ell}$ remains constant at low $\ell_{S}$. This effect is more evident when the multipole axis is plotted logarithmically, as did in this article.
@@
In summary, the mechanism generating primordial perturbations produces scalar, vector, and tensor modes. While vector modes and tensors decay with the expansion of the Universe, their contribution to the low $\ell$ signal, especially gravity waves, can be challenging to distinguish from other effects. However, polarization information can help differentiate tensor modes [27].
@@ 1.2 The acoustic peak region ($90<\ell<900)$
The intricate patterns observed in the anisotropy spectrum on intermediate scales can be attributed to gravity–driven acoustic oscillations that occurred prior to the neutralization of atoms in the universe. The frozen–in phases of these sound waves leave an imprint on cosmological parameters, lending significant constraining power to the anisotropies in the CMB.
During the era when the proton–electron plasma was closely coupled to photons, both components behaved as a unified ‘photon–baryon fluid’, wherein photons contributed primarily to pressure, and baryons imparted inertia. Gravitational potential perturbations, dominated by dark matter, drove oscillations in this fluid, with photon pressure acting as the restoring force. Despite their small amplitude ($\mathcal{O}$($10^{-5}$)), these perturbations evolved linearly, treating each Fourier mode independently as a driven harmonic oscillator. The frequency of these oscillations was determined by the sound speed in the fluid, resulting in a fluid density oscillation with a velocity out of phase and reduced amplitude.
Following the recombination of baryons and the decoupling of radiation, allowing photons to travel freely, the oscillation phases were frozen–in. Projected onto the sky, these frozen–in phases manifested as a harmonic series of peaks. The primary peak represented the mode that completed $\nicefrac{{1}}{{4}}$ of a period, reaching maximal compression. Even peaks denoted maximal under–densities, typically of smaller amplitude due to the rebound having to overcome baryon inertia. Troughs, not reaching zero power, were partially filled as they corresponded to velocity maxima [27].
@@ 1.3 The silk damping region ($\ell>900)$
The last scattering surface acquires a non-instantaneous thickness due to the recombination process, resulting in a damping effect on anisotropies at higher l's. This damping is particularly noticeable on scales smaller than the thickness of the last scattering surface. Another perspective is to view the photon–baryon fluid as having imperfect coupling, leading to diffusion between the two components and a gradual reduction in the amplitudes of oscillations over time. These effects collectively contribute to a phenomenon known as Silk damping, effectively truncating the anisotropies at multipoles exceeding approximately 2000 [27].
The paper is organized as follows: in Section 2 we define the slow–roll parameters and we describe how to calculate the cosmological parameters into the slow–roll approximation. Sections $3$, $4$, and $5$ are devoted to calculating the CMB angular power spectrum obtained from CAMB for the Starobinsky inflationary model, the generalized Starobinsky inflationary model, and the chaotic inflationary model with a step. Finally, the conclusions are discussed in Section $6$.
@@ 2 Slow–roll approximation and cosmological parameters
The slow–roll approximation is a standard technique in inflationary cosmology to solve the movement equations and the scalar and tensor equation of perturbations. This approximation is characterized by a set of parameters that establish when the approximation is valid or not and are given in terms of the inflationary potential and its derivatives in the following way [28, 29]
@@
where $\epsilon_{\nu}$, $\eta_{\nu}$, $\xi_{\nu}^{2}$ and $\omega_{\nu}^{3}$ are called slow–roll parameters, while $V$ is the potential.
In order to compute the CMB power spectrum, we need to calculate the cosmological quantities in terms of the slow–roll parameters as inputs in the CAMB cod
@@
We also need as input the value of the amplitude of the scalar and tensor power spectra, for which we use the value calculated numerically [30].
@@ 3 The Starobinsky inflationary model
where $M=1.13\times 10^{-5}$ [31] to fit the amplitude of the scalar power spectrum, and $\phi$ is the inflaton.
From Eqs. (4)–(7) with the potential given by Eq. (14) we calculate analytically the the slow–roll parameters for the Starobinsky inflationary model, and following the notation of Renzi et al. [16] we found that they are given by:
@@
where $F_{1}=e^{\sqrt{\frac{2}{3}}\phi}$.
@@
In Fig. 1(b) is observed the large angular scales ($\ell\lesssim 90$) and its relative error is shown in Fig. 2(b). At these scales the contribution to CMB anisotropies are from perturbations that were superhorizon at recombination [32]. In Fig. 1(c) is observed the intermediate angular scales ($90<\ell\lesssim 900$) and its relative error is shown in Fig. 2(c). In Fig. 1(d) is observed the large angular scales ($\ell>900$) and its relative error is shown in Fig. 2(d).
@@
We can easily observe in the relative error plots that the CAMB–reproduced model aligns well with Planck $2018$ observations, as indicated by the small errors. However, at smaller angular scales, results exhibit more variability, as evident from the oscillatory behavior of the errors in this section. To assess the overall balance of these errors, let’s examine the L2222 norm.
@@ 3.1 The Acoustic Peaks
In Table 1, the acoustic peaks found in the intermediate scale of the spectrum are listed, along with the troughs in this region. We observed that the positions on the multipole axis of the peaks are close to each other for both observed and computed values. Considering the errors of the observed values, all computed peaks, except for the third one, are in agreement with the Planck satellite. The percentage errors of the calculated multipole for peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $0.18$, $1.27$, $0.20$, $0.22$, and $0.52$, respectively. However, the differences between the observed and computed amplitudes of the peaks and troughs appear to be greater. The percentage errors of the calculated amplitudes for peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $4.84$, $10.86$, $3.90$, $5.18$, and $1.89$, respectively. These errors indicate that, indeed, there is less agreement with the observations in the amplitude of the acoustic peaks than with their multipole.
@@
3.2 Cosmic parameters
In Table 2, some cosmic parameters are calculated theoretically from the Starobinsky inflationary model are compared with their respective observational values from the Planck satellite. Considering the uncertainties, all the parameters are in agreement between their theoretical and observational values.
Cosmic Paramenter Symbol Starobinsky Planck
@@ 3.2 Cosmic parameters
Table 2: Cosmic parameters obtained from the Starobinsky inflationary model compared with the cosmic parameters reported by Planck $2018$ results.
The percentage errors for the age of the universe, matter density, baryon density, dark energy density, and scalar spectral index are $0.007$, $0.16$, $0.04$, $0.09$, and $0.2$, respectively. The low percentage errors demonstrate the goodness, accuracy, and precision that the Starobinsky inflationary model exhibits when predicting the observed cosmic parameters.

### v4

@@ Abstract
This paper presents the CMB angular power spectrum obtained using the CAMB code for three different models of inflation: the Starobinsky inflationary model, the generalized Starobinsky inflationary model, and the chaotic inflationary model with a step. The results are compared with the most recent data reported for the Planck mission. An analysis of the large ($\ell\lesssim 90$), intermediate ($90\lesssim\ell\lesssim 900$), and small ($\ell\gtrsim 900$) angular scales is performed. We report the position of the peaks in the intermediate region so as the cosmological parameters obtained in each of the models: age of the universe, $\Omega_{m}$, $\Omega_{b}$, $\Omega_{\Lambda}$, $\Omega_{K}$ and $n_{\mathrm{S}}$. We also perform a Bayesian analysis using the Cobaya code to evaluate our three best-fitting models. Additionally, we generated contour plots $(n_{\mathrm{S}},r)$ for our inflationary models, taking into account the number of e–folds between the end of inflation and the completion of reheating.
@@
CMB angular power spectrum , CAMB code , Cobaya code , Reheating , Starobinsky inflationary model , generalized Starobinsky inflationary model , chaotic inflationary model with a step.
@@
[…]
[label1]organization=Yachay TECH University,School of Physical Sciences and Nanotechnology, addressline=Hda. San Jose S/N y Proyecto Yachay, city=Urcuqui, postcode=100119, state=Imbabura, country=Ecuador label3label3footnotetext:
[…]
[label2]organization=Instituto de Radioastronomia y Astrofisica, UNAM, addressline=Campus Morelia, AP 3-72, city=Mexico, postcode=58089, country=Mexico
@@ 1 Introduction
The Big Bang theory describes the development of the Universe from a much hotter and denser point to the present. It is supported by solid observational evidence: the expansion of the Universe, the prediction of light element abundances, and the existence of the Cosmic Microwave Background (CMB) radiation. The CMB is a picture of the recombination epoch ($370,000$ years after the big bang). The first prediction of the CMB was made by Alpher and Herman in $1948$ [1], and was discovered in $1965$ by Penzias and Wilson [2]. It became one of the most important observational probes of the Big–bang theory. The CMB is an almost uniform and isotropic radiation field and shows a perfect black body spectrum at a temperature of $2.72$K [3]. The CMB has fluctuations in a part of $10^{5}$ that gives rise to the structure of the present Universe [4].
Although Big Bang theory stands as a robust and widely accepted explanation for the beginning of our universe, this theory has three problems related to its initial conditions: the flatness problem, the horizon problem, and the monopole problem. To address these concerns, Alan Guth introduced the inflationary theory in $1981$ [5]. This theory proposes an epoch of accelerated expansion preceding the hot stage mentioned in the traditional Big Bang model. The simplest way to produce such an epoch is via the potential energy of a scalar field, and the field $\phi$ is called the inflaton. There are several inflationary models with their own scalar potentials, and over the years studies have been made on which of them is favored by satellite data [6, 7, 8, 9]. One notable outcome of inflation is that it provides a mechanism to generate the primordial scalar perturbations and produce an almost scale-invariant scalar power spectrum.
@@
Using the CAMB code we study for each model the three regions of the CMB angular power spectrum: a) the Sachs–Wolfe plateau region ($\ell<90$) which correspond to large angular scales ($\theta>2^{\circ}$), b) the acoustic peak region ($90090<\ell<$ 900), and c) the silk damping region ($\ell>900$). It is important to note that Macorra et al. have used the CAMB code with full COSMO--MC (MCMC) calculation to give a solution to one of the most recent conflicts in Modern Cosmology: the Hubble tension [23]. Also, Purkayastha et al. have estimate CMB E mode signal over large angular scales [24]. Additionally, we have used the Cobaya code [25] in order to perform a Bayesian analysis of our three best inflationary models.
To study CMB, we utilize the radiation angular power spectrum, which is measurable and theoretically predictable. The radiation angular power spectrum $C_{\ell}$ is defined by:
@@
$C_{\ell}$ depends only on $\ell$ since the statistical properties are required to be independent of the choice of the origin of $\theta-\phi$, that is, rotational invariance, or from a physical point of view, isotropy [26]. The larger is $\ell$, the smaller is the angular scale at which the spherical harmonics have variation [27].
Also (and this is the case for this article), the $C_{\ell}$ spectrum is conventionally plotted as $\ell(\ell+1)C_{\ell}\overline{T}^{2}/{2\pi}$, measuring the power per logarithmic interval in $\ell$. In such plot, a scale–invariant spectrum looks horizontal [26].
@@ 1.1 The Sachs-–Wolfe plateau region ($\ell$ $<$ 90)
Assuming a nearly scale–invariant spectrum of density perturbations in early times ($n_{S}\approx 1$), meaning gravitational potential fluctuations that are independent of $k$, then $\ell(\ell+1)C_{\ell}$ remains constant at low $\ell_{S}$. This effect is more evident when the multipole axis is plotted logarithmically, as we did in this article.
@@
In summary, the mechanism generating primordial perturbations produces scalar, vector, and tensor modes. While vector modes and tensors decay with the expansion of the Universe, their contribution to the low $\ell$ signal, especially gravity waves, can be challenging to distinguish from other effects. However, polarization information can help to differentiate tensor modes [28].
@@ 1.2 The acoustic peak region ($90<\ell<900)$
The intricate patterns observed in the anisotropy spectrum on intermediate scales can be attributed to gravity–driven acoustic oscillations that occurred prior to the neutralization of atoms in the universe. The frozen–in phases of these sound waves leave an imprint on the cosmological parameters, lending significant constraining power to the anisotropies in the CMB.
During the era when the proton–electron plasma was closely coupled to photons, both components behaved as a unified ‘photon–baryon fluid’, wherein photons contributed primarily to pressure and baryons imparted inertia. Gravitational potential perturbations, dominated by dark matter, drove oscillations in this fluid, with photon pressure acting as the restoring force. Despite their small amplitude ($\mathcal{O}$($10^{-5}$)), these perturbations evolved linearly, treating each Fourier mode independently as a driven harmonic oscillator. The frequency of these oscillations was determined by the sound speed in the fluid, resulting in a fluid density oscillation with a velocity out of phase and a reduced amplitude.
Following the recombination of baryons and the decoupling of radiation, allowing photons to travel freely, the oscillation phases were frozen–in. Projected onto the sky, these frozen–in phases manifested as a harmonic series of peaks. The primary peak represented the mode that completed $\nicefrac{{1}}{{4}}$ of a period, reaching maximum compression. Even peaks denoted maximal under–densities, typically of smaller amplitude due to the rebound having to overcome baryon inertia. The troughs, which did not reach zero power, were partially filled as they corresponded to the maximum velocity [28].
@@ 1.3 The silk damping region ($\ell>900)$
The last scattering surface acquires a non–instantaneous thickness due to the recombination process, resulting in a damping effect on anisotropies at higher l's. This damping is particularly noticeable on scales smaller than the thickness of the last scattering surface. Another perspective is to view the photon–baryon fluid as having imperfect coupling, leading to diffusion between the two components and a gradual reduction in the amplitudes of oscillations over time. These effects collectively contribute to a phenomenon known as Silk Damping, effectively truncating the anisotropies at multipoles exceeding approximately 2000 [28].
The paper is organized as follows: in Section 2 we define the slow–roll parameters and we describe how to calculate the cosmological parameters into the slow–roll approximation. Sections $3$, $4$, and $5$ are devoted to calculating the CMB angular power spectrum obtained from CAMB for the Starobinsky inflationary model, the generalized Starobinsky inflationary model, and the chaotic inflationary model with a step. In Section $6$ we made a comparison of our three best fitting models through the calculation of $\chi^{2}$. Section $7$ shows the results of the Bayesian Analysis with the variation of the effective number of neutrino species $\Delta N_{\text{eff}}$ in order to determine which of our three best fitting models is most favored. Section 8 focuses on the incorporation of reheating analysis and its predictions in the $(n_{\mathrm{S}},r)$ contour plots for all of our inflationary models. Finally, the conclusions are discussed in Section $9$.
@@ 2 Slow–roll approximation and cosmological parameters
The slow–roll approximation is a standard technique in inflationary cosmology to solve the movement equations and the scalar and tensor equation of perturbations. This approximation is characterized by a set of parameters that establish when the approximation is valid or not and are given in terms of the inflationary potential and its derivatives in the following way [29, 30]:
@@
where $\epsilon_{\nu}$, $\eta_{\nu}$, $\xi_{\nu}^{2}$, and $\omega_{\nu}^{3}$ are known as the slow–roll parameters, while $V$ represents the potential characterizing the inflationary model under consideration. The prime denotes a derivative with respect to the scalar field $\phi$.
In order to compute the CMB power spectrum we need to configure each .ini file for the CAMB code, it is essential to determine the initial value of the scalar field at the beginning of inflation $\phi_{\text{ini}}$ [30]. This value can be computed using the expression for the number of e-folds, $N_{e}$, under the slow–roll approximation, as follows:
$N_{e}\simeq\int_{\phi_{\text{end}}}^{\phi}\dfrac{V}{V^{\prime}}\mathrm{d}\phi,$ (8)
where $\phi_{\text{end}}$ is the value of the scalar field at the end of inflation, determined by the condition $\epsilon_{\nu}=1$, where $\epsilon_{\nu}$ is one of the slow–roll parameters defined in Eq. (4). The number of e–folds, $N_{e}$, is fixed at 60 [30].
Using the value of $\phi_{\text{ini}}$, we calculate the cosmological quantities in terms of the slow–roll parameters as,
@@
In addition to the initial conditions for the scalar field, we also require the amplitudes of the scalar power spectrum as input for the CAMB code. To perform the numerical calculation of the scalar power spectrum, the first step involves solving the background equations. With the background solutions $a(t)$ and $\phi(t)$ established, we construct the equations for the scalar perturbations, allowing us to numerically calculate the scalar power spectrum [14, 31].
@@ 3 The Starobinsky inflationary model
where $M=1.30\times 10^{-5}$ [32] to fit the amplitude of the scalar power spectrum, and $\phi$ is the inflaton.
From Eqs. (4)–(7) with the potential given by Eq. (15) we calculate analytically the slow–roll parameters for the Starobinsky inflationary model, and following the notation of Renzi et al. [16] we found that they are given by:
@@
where $F_{1}=e^{\sqrt{\frac{2}{3}}\phi}$. Using Eq. (8) we obtain the expression for $\phi_{\text{ini}}$,
$\phi_{\text{ini}}=\dfrac{1}{6}\left[-3\sqrt{6}e^{\sqrt{\frac{2}{3}}\phi_{\text{end}}}-4\sqrt{6}N_{e}+6\phi_{\text{end}}-3\sqrt{6}W_{-1}\left(-e^{-e^{\sqrt{\frac{2}{3}}\phi_{\text{end}}}-\frac{4Ne}{3}+\sqrt{\frac{2}{3}}\phi_{\text{end}}}\right)\right],$ (20)
where $W_{-1}$ is the Lambert function, and $\phi_{\text{end}}$ is obtained from Eq. (16) for making $\epsilon_{\nu}=1$,
$\phi_{\text{end}}=\sqrt{\dfrac{3}{2}}\ln\left[\dfrac{1}{3}\left(3+2\sqrt{3}\right)\right].$ (21)
In Fig. 1, we show the evolution of the scalar field $\phi(t)$ as a function of time $t$ for the Starobinsky inflationary model. The figure illustrates the oscillation of the scalar field $\phi(t)$ after the end of inflation.
Figure 1: Evolution of the scalar field $\phi(t)$ as a function of time $t$ of the Starobinsky inflationary model.
Using Eq. (14) we obtain that the value for the scalar–to–tensor ratio $r$ for the Starobinsky inflationary model is,
$r=0.0029639.$ (22)
Figs. 2(a) and 2(b) shows that the slow–roll parameters (16) and (17) for the Starobinsky inflationary model are satisfied.
Figure 2: Behaviour of the slow–roll parameters $\epsilon_{\nu}$ and $\eta_{\nu}$ with respect to the number of e–folds $N_{e}$, described by the Starobinsky inflationary model.
@@
In Fig. 3(b) is observed the large angular scales ($\ell\lesssim 90$) and its relative error is shown in Fig. 4(b). At these scales the contribution to CMB anisotropies is from perturbations that were superhorizon at recombination [33]. In Fig. 3(c) is observed the intermediate angular scales ($90<\ell\lesssim 900$) and its relative error is shown in Fig. 4(c). In Fig. 3(d) is observed the large angular scales ($\ell>900$) and its relative error is shown in Fig. 4(d).
@@
We can easily observe in the relative error plots that the CAMB–reproduced model aligns well with Planck $2018$ observations, as indicated by the small errors. However, at smaller angular scales, results exhibit more variability, as evident from the oscillatory behavior of the errors in this section. To assess the overall balance of these errors, let us examine the L2222 norm.
@@ 3.1 The Acoustic Peaks
In Table 1, the acoustic peaks found in the intermediate scale of the spectrum are listed, along with the troughs in this region. We observed that the positions on the multipole axis of the peaks are close to each other for both observed and computed values. Considering the errors of the observed values, all computed peaks, except the third one, agree with the Planck satellite. The percentage errors of the calculated multipole for peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $0.18$, $1.27$, $0.20$, $0.22$, and $0.52$, respectively. However, the differences between the observed and computed amplitudes of the peaks and troughs appear to be greater. The percentage errors of the calculated amplitudes for peak $1$, trough $1$, peak $2$, trough $2$, and peak $3$ are $4.84$, $10.86$, $3.90$, $5.18$, and $1.89$, respectively. These errors indicate that, indeed, there is less agreement with the observations in the amplitude of the acoustic peaks than with their multipole.
@@
3.2 Cosmological parameters
In Table 2, some cosmological parameters theoretically calculated from the Starobinsky inflationary model are compared with their respective observational values from the Planck satellite. Considering the uncertainties, all the parameters are in agreement between their theoretical and observational values.
Cosmological Parameter Symbol Starobinsky Planck
@@ 3.2 Cosmological parameters
Table 2: Cosmological parameters obtained from the Starobinsky inflationary model compared with the cosmological parameters reported by Planck $2018$ results.
The percentage errors for the age of the universe, matter density, baryon density, dark energy density, and scalar spectral index are $0.007$, $0.16$, $0.04$, $0.09$, and $0.2$, respectively. The low percentage errors demonstrate the goodness, accuracy, and precision that the Starobinsky inflationary model exhibits when predicting the observed cosmological parameters.